What is the vertex form of the equation?
y = x^2 + 4x – 3



Answer by ewatrrr(24785)

 
Hi,
What is the vertex form of the equation?
y = x^2 + 4x – 3
y = x^2 + 4x +4 -4 – 3
y=(x+2)^2 - 7 Vertex form: V(-2,-7)